# Cwikel–Lieb–Rozenblum type inequalities for Hardy–Schrödinger operator

### v1

@@ 1. Introduction and Main Results
where $V_{+}(x)=\max(V(x),0)$. In particular, since $N(0,-\Delta-V)$ is always an integer, (1) implies that $N(0,-\Delta-V)=0$ if $\|V_{+}\|_{L^{d/2}(\mathbb{R}^{d})}$ is small enough, which can be deduced from Sobolev’s inequality
@@
$N(0,-\Delta-\lambda V)$ $=\frac{1}{(2\pi)^{d}}|\{(p,x)\in\mathbb{R}^{d}\times\mathbb{R}^{d}:|p|^{2}-\lambda V(x)<0\}+o(\lambda^{d/2})_{\lambda\to\infty}$
@@
On the other hand, it is well-known that (6) does not hold for $\gamma=0$ [EF06]. In fact, even the corresponding Sobolev inequality does not hold with $\|u\|_{\dot{H}^{1}}^{2}$ replaced by $\langle u,\mathcal{L}u\rangle$. To be precise, it was proved by Filippas–Tertikas in [FT02] (see also Musina’s remarks in [M09]) that
@@
where the power of the logarithmic weight is optimal. Here $B$ denotes the unit ball in $\mathbb{R}^{d}$. By a duality argument similar to (3), the Hardy–Sobolev inequality (7) is equivalent to the fact that $\mathcal{L}-W\geq 0$ on $L^{2}(B)$, as quadratic forms with Dirichlet boundary conditions, if
@@
Our first new result is an extension of the above Hardy–Sobolev inequality (7) concerning the number of negative eigenvalues of $\mathcal{L}-W$ on $L^{2}(\mathbb{R}^{d})$.
@@
Our proof of Lemma 4 uses an analogue of Rumin’s method [R11], plus the precise spectral property of $\mathcal{L}_{B}$ which has been understood by Vazquez–ZuaZua [VZ00]. Although the result there holds only the the unit ball, its application to the whole $\mathbb{R}^{d}$ is made possible due to the relation
@@
In fact, the proof of Theorem 6 is also valid when $s=1$ and $d\geq 3$ and, in fact, simplifies at some points. Thus, we obtain an alternative proof of Theorem 1. In the local case, however, we feel that the first proof is more direct, which motivated us to present it first.
@@ 2.1. Improved Hardy Inequalities
The above estimate comes from the fact that the lowest nontrivial eigenvalue of the Laplace–Beltrami operator on the unit sphere $\mathbb{S}^{d-1}$ is equal to $d-1$. On the radial part, we recall the following result from Musina [M09].
@@ 2.2. Strauss type Estimate
$\leq\int_{B}\left|\sum_{k\geq 1}\frac{1}{\sqrt{\lambda_{0,k}}}\overline{\varphi_{0,k}(y)}\varphi_{0,k}(x)\,{\rm d}y\right|^{2}\leq\sum_{k\geq 1}\frac{|\varphi_{0,k}(x)|^{2}}{\lambda_{0,k}}.$
@@ 2.3. Conclusion of Theorem 1
$\geq\frac{1}{4}\int_{B}\frac{\rho(x)}{|x|^{2}|\ln|x||^{2}}{\rm d}x\gtrsim_{d}\int_{B}\frac{\rho(x)^{\frac{d}{d-2}}}{(1+|\ln|x|)^{1+\frac{d}{d-2}}}{\rm d}x$
@@
$\int_{1}^{\infty}w(r)u(r)^{2}r^{d-1}\,{\rm d}r=\int_{0}^{1}\tilde{w}(s)\tilde{u}(s)^{2}s^{d-1}\,{\rm d}s\,.$
@@
This, together with (18) and (19) completes the proof of Theorem 1. ∎
@@ 3.2. Equivalence of Sobolev and CLR Inequalities
    there is a measurable, a.e. positive function $\omega$ such that, if $u\in\operatorname{dom}t$ is non-negative then $\min(u,\omega)\in\operatorname{dom}t$ and $t[\min(u,\omega)]\leq t[u]$. Moreover, there is a form core $\mathcal{Q}$ of $t$ such that $\omega^{-1}\mathcal{Q}$ is dense in $L^{2}(X,\omega^{2\kappa/(\kappa-1)}{\rm d}x)$.
@@
This is the assertion (ii) that we wanted to prove under the assumption $w\in L^{\infty}(X)$. Now we consider the general case, where $w$ is finite a.e. As we mentioned before, we may also assume that $w$ is bound away from zero. The Hilbert space $\mathfrak{H}$ is defined as before, but now we only know that $L^{2}(X)\subset\mathfrak{H}$. Given $\delta>0$ we consider the quadratic form $v\mapsto t_{\delta}[v]:=t[v]+\delta\int_{X}|v|^{2}\,{\rm d}x$ in the Hilbert space $\mathfrak{H}$ with form domain $\operatorname{dom}t$. Let us show that $t_{\delta}$ is closed in $\mathfrak{H}$. Thus, let $(v_{j})\subset\operatorname{dom}t$ and $v\in\mathfrak{H}$ such that $v_{j}\to v$ in $\mathfrak{H}$ and such that $(v_{j})$ is Cauchy with respect to $t_{\delta}$. Since $t$ is nonnegative, we infer that $(v_{j})$ is Cauchy with respect to the norm in $L^{2}(X)$ and hence convergent in $L^{2}(X)$. Passing to a.e.-convergent subsequences and recalling that $w$ is finite a.e., it is easy to see that $v\in L^{2}(X)$ and that $v_{j}\to v$ in $L^{2}(X)$. Therefore the assumed closedness of $t$ implies that $v\in\operatorname{dom}t$ and that $t[v_{j}-v]\to 0$. Thus, we have shown that $t_{\delta}$ is closed in $\mathfrak{H}$. Let us denote by $A_{\delta}$ the nonnegative operator in $\mathfrak{H}$ generated by $t_{\delta}$. The form $t_{\delta}$ satisfies Assumptions 13 and we deduce, as before, that
@@
We apply Theorem 14 in the measure space $X=B$ with Lebesgue measure and with $t[u]=\|\mathcal{L}_{s}^{1/2}u\|^{2}$. Items (a) and (b) in Assumption 13 are clearly satisfied. The first part of item (c) is satisfied with $\omega(x)=|x|^{-(d-2s)/2}$, as follows from the ground state substitution formula in the [FLS09] on Hardy–Lieb–Thirring. The second part is satisfied, for any $1<\kappa<\infty$, since $C^{\infty}_{c}(B\setminus\{0\})$ is a form core of $\mathcal{L}_{s}$. Thanks to Lemma 12, the Sobolev inequality (24) is satisfied with $q=\frac{2d}{d-2s}$ and $w=(1+|\ln|x||)^{\frac{d-s}{s}}$. The assertion follows now from the CLR inequality (25). ∎
@@
thus proving the claimed identity (27). Identity (27) is the analogue of Lemma 10. As in the proof of Corollary 11, we deduce that
@@ 3.3. Localization
This formula is due to Michael Loss and appeared in [LW88]. By the triangle inequality we have the pointwise estimate
@@ Acknowledgements
Partial support through the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Germany’s Excellence Strategy EXC - 2111 - 390814868 and through TRR 352 – Project-ID 470903074 (G.K. Duong, R.L. Frank, P.T. Nam), through the U.S. National Science Foundation grant DMS-1954995 (R.L. Frank), and through the Czech Science Foundation Project GA22-17403S (T.M.T. Le, P.T. Nam, P.T. Nguyen) is acknowledged.

### v2

@@ 1. Introduction and Main Results
where $V(x)_{+}=\max(V(x),0)$. Here the notation $\lesssim_{d}$ means that the implicit constant on the right hand side depends only on the dimension $d$. In particular, since $N(0,-\Delta-V)$ is always an integer, (1) implies that $N(0,-\Delta-V)=0$ if $\|V_{+}\|_{L^{d/2}(\mathbb{R}^{d})}$ is small enough, which can be deduced from Sobolev’s inequality
@@
$N(0,-\Delta-\lambda V)$ $=\frac{1}{(2\pi)^{d}}|\{(p,x)\in\mathbb{R}^{d}\times\mathbb{R}^{d}:|p|^{2}-\lambda V(x)<0\}|+o(\lambda^{d/2})_{\lambda\to\infty}$
@@
On the other hand, it is well-known that (6) does not hold for $\gamma=0$ [EF06]. In fact, even the corresponding Sobolev inequality does not hold with $\|u\|_{\dot{H}^{1}}^{2}$ replaced by $\langle u,\mathcal{L}u\rangle$. However, there is a remarkable replacement for the Sobolev inequality in the restricted case where $\mathbb{R}^{d}$ is replaced by the unit ball $B$. To be precise, it was proved by Filippas–Tertikas in [FT02] (see also Musina’s remarks in [M09]) that
@@
where the power of the logarithmic weight is optimal. By a duality argument similar to (3), the Hardy–Sobolev inequality (7) is equivalent to the fact that $\mathcal{L}-W\geq 0$ on $L^{2}(B)$, as quadratic forms with Dirichlet boundary conditions, if
@@
Our first new result is an extension of the above Hardy–Sobolev inequality concerning the number of negative eigenvalues of $\mathcal{L}-W$ on $L^{2}(\mathbb{R}^{d})$. In particular, it allows to extend (7) to the whole $\mathbb{R}^{d}$.
@@
Our proof of Lemma 4 uses an analogue of Rumin’s method [R11], plus the precise spectral property of $\mathcal{L}_{B}$ which has been understood by Vázquez–ZuaZua [VZ00]. Although the result there holds only the unit ball, its application to the whole $\mathbb{R}^{d}$ is made possible due to the relation
@@
In fact, the proof of Theorem 6 is also valid when $s=1$ and $d\geq 3$ and simplifies at some points. Thus, we obtain an alternative proof of Theorem 1. In the local case, however, we feel that the first proof is more direct, which motivated us to present it first.
@@ 2.1. Improved Hardy Inequalities
The above estimate comes from the fact that the lowest nontrivial eigenvalue of the Laplace–Beltrami operator on the unit sphere $\mathbb{S}^{d-1}$ is equal to $d-1$.
On the radial part, we recall the following result from Musina [M09].
@@ 2.2. Strauss type Estimate
$\leq\int_{B}\left|\sum_{k\geq 1}\frac{1}{\sqrt{\lambda_{0,k}}}\overline{\varphi_{0,k}(y)}\varphi_{0,k}(x)\right|^{2}\,{\rm d}y\leq\sum_{k\geq 1}\frac{|\varphi_{0,k}(x)|^{2}}{\lambda_{0,k}}.$
@@ 2.3. Conclusion of Theorem 1
$\geq\frac{1}{4}\int_{B}\frac{\rho(x)}{|x|^{2}|\ln|x||^{2}}{\rm d}x\gtrsim_{d}\int_{B}\frac{\rho(x)^{\frac{d}{d-2}}}{(1+|\ln|x||)^{1+\frac{d}{d-2}}}{\rm d}x$
@@
$\int_{1}^{\infty}w(r)u(r)^{2}r^{d-1}\,{\rm d}r=\int_{0}^{1}\tilde{w}(s)\tilde{u}(s)^{2}s^{d-1}\,{\rm d}s,$
@@
This, together with (18) and (19), completes the proof of Theorem 1. ∎
@@ 3.2. Equivalence of Sobolev and CLR Inequalities
    there is a measurable, a.e. positive function $\mu$ such that, if $u\in\operatorname{dom}t$ is non-negative then $\min(u,\mu)\in\operatorname{dom}t$ and $t[\min(u,\mu)]\leq t[u]$. Moreover, there is a form core $\mathcal{Q}$ of $t$ such that $\mu^{-1}\mathcal{Q}$ is dense in $L^{2}(X,\mu^{2\kappa/(\kappa-1)}{\rm d}x)$.
@@
This theorem for $w=1$ appears in the paper [FLS09]. It is based on a method due to Li and Yau [LY83] with an improvement in [BSR87] and generalizes a theorem of Levin and Solomyak [LS97].
@@
Thus, assume that $w$ is bounded away from zero. To better explain the strategy of the proof, let us first assume, in addition, that $w\in L^{\infty}(X)$. Then the Hilbert space $\mathfrak{H}:=L^{2}(X,w^{-(q-2)/2}{\rm d}x)$ coincides, with equivalent norm, with the Hilbert space $L^{2}(X)$. We consider the quadratic form $t$ in the Hilbert space $\mathfrak{H}$. It is clearly nonnegative and, by our assumptions on $w$, it is closed. (It is in the proof of closedness that we use the assumption $w\in L^{\infty}(X)$.) Thus it generates a nonnegative operator $A$ in $\mathfrak{H}$. Moreover, it satisfies Assumption 13, since we are assuming that the corresponding assumption is satisfied for the original form $t$ in the Hilbert space $L^{2}(X)$. Applying [FLS09, Theorem 2.1] to the operator $A$, we obtain that for any $0\leq U\in L^{\kappa}(X,w^{-(q-2)/2}{\rm d}x)$ one has
@@
This is the assertion (ii) that we wanted to prove under the assumption $w\in L^{\infty}(X)$.
Now we consider the general case, where $w$ is finite a.e. As we mentioned before, we may also assume that $w$ is bound away from zero. The Hilbert space $\mathfrak{H}$ is defined as before, but now we only know that $L^{2}(X)\subset\mathfrak{H}$. Given $\delta>0$ we consider the quadratic form $v\mapsto t_{\delta}[v]:=t[v]+\delta\int_{X}|v|^{2}\,{\rm d}x$ in the Hilbert space $\mathfrak{H}$ with form domain $\operatorname{dom}t$. Let us show that $t_{\delta}$ is closed in $\mathfrak{H}$. Thus, let $(v_{j})\subset\operatorname{dom}t$ and $v\in\mathfrak{H}$ such that $v_{j}\to v$ in $\mathfrak{H}$ and such that $(v_{j})$ is Cauchy with respect to $t_{\delta}$. Since $t$ is nonnegative, we infer that $(v_{j})$ is Cauchy with respect to the norm in $L^{2}(X)$ and hence convergent in $L^{2}(X)$. Passing to a.e.-convergent subsequences and recalling that $w$ is finite a.e., it is easy to see that $v\in L^{2}(X)$ and that $v_{j}\to v$ in $L^{2}(X)$. Therefore the assumed closedness of $t$ implies that $v\in\operatorname{dom}t$ and that $t[v_{j}-v]\to 0$. Thus, we have shown that $t_{\delta}$ is closed in $\mathfrak{H}$.
Let us denote by $A_{\delta}$ the nonnegative operator in $\mathfrak{H}$ generated by $t_{\delta}$. The form $t_{\delta}$ satisfies Assumptions 13 and we deduce, as before, that
@@
We apply Theorem 14 in the measure space $X=B$ with Lebesgue measure and with $t[u]=\|\mathcal{L}_{s}^{1/2}u\|^{2}$. Items (a) and (b) in Assumption 13 are clearly satisfied. The first part of item (c) is satisfied with $\mu(x)=|x|^{-(d-2s)/2}$, as follows from the ground state substitution formula in the [FLS09] on Hardy–Lieb–Thirring. The second part is satisfied, for any $1<\kappa<\infty$, since $C^{\infty}_{c}(B\setminus\{0\})$ is a form core of $\mathcal{L}_{s}$. Thanks to Lemma 12, the Sobolev inequality (24) is satisfied with $q=\frac{2d}{d-2s}$ and $w=(1+|\ln|x||)^{\frac{d-s}{s}}$. The assertion follows now from the CLR inequality (25). ∎
@@
thus proving the claimed identity (27).
Identity (27) is the analogue of Lemma 10. As in the proof of Corollary 11, we deduce that
@@ 3.3. Localization
This formula is due to Michael Loss and appeared in [LW88].
By the triangle inequality we have the pointwise estimate
@@ Acknowledgements
We thank the referees for helpful suggestions. Partial support through the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Germany’s Excellence Strategy EXC - 2111 - 390814868 and through TRR 352 – Project-ID 470903074 (G.K. Duong, R.L. Frank, P.T. Nam), through the U.S. National Science Foundation grant DMS-1954995 (R.L. Frank), and through the Czech Science Foundation Project GA22-17403S (T.M.T. Le, P.T. Nam, P.T. Nguyen) is acknowledged.